# High Resolution UDF Meshing via Iterative Networks

### v2

@@ 1 Introduction
Implicit Neural Representations [Sitzmann20, Gropp20, Novello22] have become powerful tools to accurately model objects whose topology is a priori unknown and without being bound to a specific mesh resolution. Those based on Occupancy Fields [Mescheder19] or Signed Distance Functions (SDFs) [Park19c] are best at handling watertight surfaces. They can be used to represent open surfaces by wrapping a thin volume around them, but this is less than ideal. Unsigned Distance Fields (UDFs) offer a more effective alternative [Chibane20b, Zhou22, Zhou24, Ren23, Zhou23, Fainstein24, Xu25] and are used by many surface reconstruction methods [Liu23a, Long22, Li25b]. In a typical scenario, a neural network is used to parametrize the field, which is then converted into an explicit mesh for downstream tasks through a meshing process. There are many algorithms to do this conversion well when the field features sign changes to signal surface locations [Lewiner03, Ju02, Shen23, Ren24]. However, there are no such sign changes in the UDF case and the field is expected to be exactly zero at the surface location.
@@
MeshUDF [Guillard22b] NSD-UDF [Stella24] DMUDF [Zhang23b] Ours - iter. 1 Ours - iter. 2 Ours - iter. 6
@@
Most current approaches to meshing UDFs [Zhou24, Zhang23b, Hou23a, Stella24] rely on Marching Cubes or Dual Contouring-like algorithms that operate on the vertices of individual voxels to recover the surface that may or may not traverse them. However neural UDFs are notoriously noisy [Guillard22b, Stella24, Hou23a], which makes it hard to retrieve the correct surface. Counterintuitively - and contrary to SDFs - increasing the meshing resolution to retrieve finer surface details actually worsens the problem, as shown in the results section (Fig. 5). This is because, at high resolutions, UDF values often become noisier or ambiguous within a voxel, and these local, single-pass methods lack the necessary context to infer the surface reliably, resulting in poorly recovered geometry. As shown in Fig. 1, when the neural UDF fails to reach sufficiently low values, a hole appears in the reconstructed surface. This is particularly apparent around the planes’s cockpit, as shown in the upper right corner of the figure where existing methods do not reliably detect a surface, even though there is a local minimum region. Thus, looking beyond single voxels is necessary, as attempted in [Chen22b, Guillard22b]. However, these attempts rely on basic propagation heuristics or a simple increase of the receptive fields and still fail at high resolutions, as also shown in the results section.
@@ 2 Related Work
Signed and unsigned distance fields are now in widespread use to represent 3D surfaces, which are typically parameterized by a latent vector that is decoded by a deep network [Mescheder19, Park19c, Chibane20b]. Shapes are not tied to a specific resolution, but only to the representational power of the deep network in use. Neural signed distance fields (SDFs) are limited to watertight shapes but are relatively easy to mesh, whereas unsigned distance fields (UDFs) can represent open surfaces. They are used in volume rendering, shape reconstruction from multi-view and shape generation with a variety of techniques [Liu23a, Long22, Li25b, DeLuigi23], but they are much harder to mesh.
@@
Marching Cubes (MC) [Lorensen87, Lewiner03] and Dual Contouring (DC) [Ju02] are widely used to mesh signed fields by detecting sign transitions within voxels of a 3D grid. MC interpolates vertex positions along cell edges, while DC places them inside cells and optimizes their locations. Deep-learning adaptations, such as Neural Marching Cubes (NMC) [Chen21c] and Neural Dual Contouring (NDC) [Chen22b], improve on the classical methods by eliminating the need for manually defined rules or accurate field gradients, respectively. FlexiCubes [Shen23] yields impressive performance on neural SDFs used for shape optimization while McGrids [Ren24] improves the scalability of surface extraction at high resolution. In all methods increasing the resolution, or the number of points used, typically produces more detailed meshes at a higher computational cost. However, as they rely on sign changes, none of these methods is directly applicable to UDFs.
@@
NDF [Chibane20b] uses a ball-pivoting algorithm to mitigate this problem. Unfortunately, its reliance on dense point clouds and expensive processing takes hours per mesh and makes it impractical. Methods such as MeshUDF [Guillard22b] and CAP-UDF [Zhou22, Zhou24] assign pseudo-signs to the field values by comparing gradients so that MC can be used, but they are subject to artifacts, especially on complex geometries. MeshUDF partially conditions its pseudo-signs on previously computed voxels by defining a specific voxel-exploration order, but it does so in a single pass and with a heuristic that targets simple shapes such as garments. DCUDF [Hou23a] extracts an $\epsilon$-level set using MC, refines it by minimizing a loss function, and cuts it to a single thin surface using a min-cut algorithm [Boykov04]. It produces smooth surfaces but requires much manual tuning, long processing times, while sometimes failing to properly cut the inflated surface. While the method refines the vertex locations with an optimization procedure, the mesh extraction is still carried out in a single pass. In NSD-UDF [Stella24], a neural network predicts pseudo-signs starting from local UDF values and gradients, which can be meshed using SDF-like triangulation methods such as MC and DualMesh-UDF [Zhang23b]. However, all these methods struggle at high resolutions on neural UDFs, where gradients and field values can become very noisy. This interferes with the proper generation of pseudo-signs, surface cuts, or dual vertices, depending on the specific method being used. This can yield inconsistencies and holes in the reconstructed surfaces, which our iterative approach is designed to prevent.
Since DC is a popular alternative to MC, it has also been adapted to work with UDFs. DualMesh-UDF [Zhang23b] prunes grid cells via a UDF-based threshold and refines vertices using a DC-like quadratic equation. While preserving sharp features, it relies on hand-crafted filtering rules that can produce holes and require manual tuning. This tuning can be partially avoided by using [Stella24] for pre-processing. However the approach still struggles at high resolutions when there is noise. Unsigned Neural Dual Contouring (UNDC) [Chen22b] uses a 3D CNN to predict dual vertex locations and connectivity. Training the network requires heavy data augmentation and ignores gradient information, limiting the method’s accuracy on neural UDFs [Zhang23b, Stella24].
@@
Recursive refinement techniques for network predictions have been demonstrated across many domains [Mnih10, Pinheiro14, Tu09, Shen17]. These approaches use the surrounding context to improve predictions [Seyedhosseini13, Durasov24a], proving especially effective for tasks such as delineation [Sironi16a, Durasov24b], human pose estimation [Newell16], semantic segmentation [Zhang18g], and depth estimation [Durasov18]. These methods have inspired us to develop the first-ever iterative approach to UDF meshing.
@@
Several approaches have been proposed to learn less noisy UDFs. Zhou et al. [Zhou22, Zhou24] enforced parallel level sets near surfaces to improve learning from sparse point clouds. Fainstein et al. [Fainstein24] use hyperbolic scaling and second-order differentiation to reduce the noise in gradients and ensure the differentiability of the field. Zhao et al. [Zhao21a] used anchor-based 3D position features to enhance UDF predictions, aiding single-view garment reconstruction. Wang et al. [Wang22b] predicted pseudo-signs from surface gradients for meshing, while Venkatesh et al. [Venkatesh20] leveraged surface orientation for better local geometry. However, precise UDF predictions remain challenging, as existing methods struggle to eliminate holes and discontinuities in meshed surfaces.
@@ 3 Method
where $\mathit{U}_{\mathcal{S}}$ is implemented by a deep network. $\mathbf{x}$ is a point in 3D space, and $d$ should be the distance from $\mathbf{x}$ to the closest points on the surface $\mathcal{S}$ corresponding to the $n$-dimensional latent vector $\mathbf{z}$ that controls the shape of $\mathcal{S}$. The difficulty with this formulation is that if $\mathit{U}_{\mathcal{S}}$ is not accurate, $\mathit{U}_{\mathcal{S}}(\mathbf{x},\mathbf{z})$ might not reach zero at places where it should, resulting in unwarranted holes in the surface reconstruction. Since unsigned fields are not differentiable everywhere, they pose a parametrization challenge, resulting in noisy representations [Guillard22b, Hou23a, Stella24]. Similarly, the gradients of $\mathit{U}_{\mathcal{S}}$, which are crucial for meshing [Guillard22b, Zhou22, Zhang23b, Stella24], can become very noisy when operating at high resolutions, also resulting in reconstruction failures.
@@ 3.1 From UDF to Triangulated Mesh
Given $\mathit{U}_{\mathcal{S}}$ of Eq. 1, we query it for UDF values and gradients on a regular grid. We then group the values in cubic cells and feed them to a neural network $f_{\theta}$, parametrized by $\theta$, whose task is to output a pseudo-sign configuration for the 8 corners of the input cell. As in [Stella24], we use a fully connected feed forward network with $2^{7}=128$ outputs, representing a one-hot encoding of the possible pseudo-sign configurations, up to a complete sign symmetry in the cell.
@@
As in NSD-UDF [Stella24], the final output of our pipeline is a pseudo-SDF, that is a UDF in which the grid cell corners have been assigned pseudo-signs so they can be triangulated using existing methods. We use Marching Cubes [Lewiner03] for its ease of use or DualMesh-UDF [Zhang23b] for its precision.
@@
Algorithms such as MeshUDF [Guillard22b] and NSD-UDF [Stella24] use manually-defined resolution-dependent thresholds that ignore cells that are too far from the surface and, thus, reduce the computational complexity of the algorithm. While this works well at lower resolutions, the high level of noise at higher resolutions makes this process error-prone, resulting in missing regions in the final mesh. Instead, we filter out cells whose vertices have a UDF value equal or above the clamping threshold of the neural UDF, set to be 0.1 in our experiments. In practice, this eliminates 85% of the cells at $256^{3}$ resolution, with no impact on the final result. For subsequent iterations, since the network outputs a probability distribution over the sign configurations, we can interpret the highest probability as the confidence of the network. We can then conservatively filter out additional cells whose confidence is extremely high (>0.999>0.999). As can be seen in Tab. 1(a), this substantially reduces the computational cost, making it comparable to that of the other baselines given in Tab. 1(b), without accuracy loss compared to meshing all cells.
Table 1: Filtering and meshing times. (a) Median L2 Chamfer Distance $\times 10^{-5}$ with 2M sample points (CD), Image Consistency (IC) and model inference times are reported at varying grid resolutions on an auto-decoder [Park19c] trained on ShapeNet [Chang15] cars. The times were measured on an NVIDIA A100 GPU. Notice that our filtering strategies put the proposed pipeline in the same range as existing methods, without impacting its accuracy.
@@
Time CAP [Zhou24] MeshUDF [Guillard22b] DCUDF [Hou23a] NSD [Stella24] UNDC [Chen22b] DMUDF [Zhang23b] Ours
@@ 4.1 Datasets and baselines
We test our method on five different 3D categories, including garments from MGN [Bhatnagar19]; cars, chairs, and planes from ShapeNet [Chang15]; and natural scenes from the 3DScene dataset [Zhou13], using four different neural UDF architectures. For garments, cars, chairs, and planes, we train a traditional auto-decoder following [Park19c, Stella24]. To showcase the capabilities of our method on downstream tasks and more precise single-shape neural architectures, we perform surface reconstruction from point clouds by training CAP-L [Zhou24] [^1] We refer to the learning architecture of CAP-UDF as CAP-L, to distinguish it from the meshing method introduced in the same paper, which we employ as a baseline. on 3D scenes and cars, and DiffUDF [Fainstein24] on cars. In the supplementary material, we provide additional results using a softplus-based auto-decoder and meshing results obtained from ground-truth UDFs. We use 300 garments and the first 20 samples each for cars, chairs, and planes. As the MGN garments are simpler and exhibit less variety than the other shapes, we use a lower resolution for them. For 3D scenes, we use three scenes from the official CAP-UDF [Zhou24] repository [^2] The repository contains five scenes, but ”Copyroom” and ”Lounge” produced bad metrics across all methods due to inconsistent normalization, so they have been excluded..
We divide the baselines into two groups, those that rely on Marching Cubes (CAP-UDF [Zhou24], MeshUDF [Guillard22b], DCUDF [Hou23a], NSD-UDF [Stella24]) and those that use Dual Contouring (UNDC [Chen22b], DualMesh-UDF [Zhang23b], NSD-UDF [Stella24]). Notably, DCUDF [Hou23a] and DualMesh-UDF [Zhang23b] have been shown to be sensitive to parameter choices. We report results obtained using both the default parameters and those that we manually tuned, denoted as “-T”. Due to implementation constraints, DualMesh-UDF and methods that depend on it use a resolution one pixel higher than the others, which we still include in the same tables. UNDC [Chen22b] failed at high resolution due to the method’s VRAM requirements. DCUDF [Hou23a] failed on certain shapes, yielding invalid metrics; these cases are excluded, making its results not directly comparable. Following official recommendations, we also evaluate DCUDF-T-nocut, a variant without the cutting step that produces double-layered meshes, on cars. All meshing methods are run without post-processing to ensure fair comparison.
@@ 4.2 Metrics
To compare extracted meshes to ground-truth surfaces, we compute three metrics: the L2 Mesh Chamfer Distance (CD) between the meshes, computed with 2M sample points as a two-way point-to-mesh distance; the Image Consistency (IC) [Guillard22b], which is the product of IoU and the cosine similarity of normal maps rendered from eight viewpoints; and the F1 score [Chen21c, Chen22b], also computed with 2M sample points, which measures the portion of the surface within an error threshold of $0.003$ from the ground-truth surface. We report the median of these metrics. It has been found to be more robust to outliers [Stella24] than the mean, which we report in the supplementary for completeness along with the standard deviation.
@@
MGN* [Bhatnagar19] ShapeNet cars [Chang15] ShapeNet chairs [Chang15] ShapeNet planes [Chang15]
@@
128 CAP-UDF [Zhou24] 16.2 69.4 79.4 53.9 53.3 83.2 378 50.9 70.3 14.0 69.5 80.6
MeshUDF [Guillard22b] 2.45 82.9 94.1 11.3 57.7 88.6 7.43 67.6 88.0 6.82 74.6 81.0
DCUDF [Hou23a] 13500 2.50 2.76 6580 7.74 14.7 17000 15.9 13.9 1880 27.6 24.5
DCUDF-T [Hou23a] 90.8 2.86 87.3 50.3 58.0 87.3 214 64.1 82.9 144 68.8 77.2
DCUDF-T-nocut [Hou23a] - - - 11.4 61.7 89.4 - - - - - -
NSD-UDF + MC [Stella24] 1.34 83.9 94.7 6.79 59.6 88.5 6.11 67.8 88.1 3.82 79.0 84.8
@@
256 CAP-UDF [Zhou24] 1.66 86.8 91.8 34.0 61.2 87.6 114 70.5 82.0 5.50 83.7 85.4
MeshUDF [Guillard22b] 0.958 89.7 95.0 13.6 62.3 88.6 27.8 72.9 87.3 3.47 85.8 85.2
DCUDF [Hou23a] 14400 4.76 3.71 346 52.6 78.2 3530 49.9 54.3 27.9 84.7 82.0
DCUDF-T [Hou23a] 4.63 86.8 95.4 347 52.6 78.2 3560 50.0 54.3 47.3 80.4 78.7
DCUDF-T-nocut [Hou23a] - - - 52.0 58.9 82.3 - - - - - -
NSD-UDF + MC [Stella24] 0.808 90.0 95.2 10.2 62.0 87.9 10.9 72.3 86.2 2.91 87.3 86.0
@@
512 CAP-UDF [Zhou24] 0.872 90.6 94.6 31.8 61.7 87.5 63.9 71.7 82.0 5.94 87.5 86.2
MeshUDF [Guillard22b] 0.798 90.6 94.8 82.7 57.0 81.7 378 61.5 65.7 12.6 88.1 84.6
DCUDF [Hou23a] 4.37 88.3 91.1 223 56.5 84.2 2950 55.0 70.1 48.7 85.5 82.2
DCUDF-T [Hou23a] 4.38 88.2 91.1 223 56.5 84.2 2000 55.0 70.1 63.0 85.4 81.3
DCUDF-T-nocut [Hou23a] - - - 43.1 60.4 85.6 - - - - - -
NSD-UDF + MC [Stella24] 0.784 90.8 94.8 56.9 58.8 83.8 295 64.7 75.7 10.0 89.4 85.1
@@
MGN* [Bhatnagar19] ShapeNet cars [Chang15] ShapeNet chairs [Chang15] ShapeNet planes [Chang15]
@@
128 UNDC [Chen22b] 1.09 87.1 94.1 13.5 61.7 86.4 29.9 69.4 81.9 2.50 82.0 86.1
DualMesh-UDF [Zhang23b] 216 68.1 68.4 952 34.4 45.5 7930 12.1 9.08 112 74.3 76.1
DualMesh-UDF-T [Zhang23b] 0.806 89.9 95.4 5.56 63.5 89.5 5.34 75.1 89.1 1.96 84.3 87.5
NSD-UDF + DualMesh-UDF [Stella24] 0.760 90.4 95.0 6.34 65.7 89.1 5.50 72.8 89.2 2.08 87.1 86.6
@@
256 UNDC [Chen22b] 0.931 89.1 91.5 82.4 52.3 71.4 293 54.2 57.8 11.6 83.6 80.7
DualMesh-UDF [Zhang23b] 176 66.3 66.4 846 34.0 45.1 8280 12.4 9.26 105 77.9 76.0
DualMesh-UDF-T [Zhang23b] 0.722 91.2 95.1 10.6 64.3 87.0 22.8 72.2 84.1 2.43 88.2 87.0
NSD-UDF + DualMesh-UDF [Stella24] 0.671 91.0 94.6 10.5 64.9 87.2 14.6 70.7 84.4 2.78 88.9 85.6
@@
512 UNDC [Chen22b] 2.39 84.8 82.8 - - - - - - - - -
DualMesh-UDF [Zhang23b] 167 63.7 63.9 870 32.5 43.0 8190 11.8 9.07 111 77.6 74.2
DualMesh-UDF-T [Zhang23b] 0.827 90.2 93.2 37.8 58.2 79.3 72.7 63.7 72.5 4.77 89.0 84.0
NSD-UDF + DualMesh-UDF [Stella24] 0.787 89.7 92.5 60.5 57.5 79.9 296 62.5 68.0 9.95 87.6 83.3
@@
CAP-L scenes [Zhou24] CAP-L cars [Zhou24] DiffUDF cars [Fainstein24]

### v3

@@ 1 Introduction
Implicit Neural Representations [35, 12, 26] have become powerful tools to accurately model objects whose topology is a priori unknown and without being bound to a specific mesh resolution. Those based on Occupancy Fields [23] or Signed Distance Functions (SDFs) [27] are best at handling watertight surfaces. They can be used to represent open surfaces by wrapping a thin volume around them, but this is less than ideal. Unsigned Distance Fields (UDFs) offer a more effective alternative [6, 44, 46, 30, 45, 11, 40] and are used by many surface reconstruction methods [20, 21, 19]. In a typical scenario, a neural network is used to parametrize the field, which is then converted into an explicit mesh for downstream tasks through a meshing process. There are many algorithms to do this conversion well when the field features sign changes to signal surface locations [18, 15, 32, 29]. However, there are no such sign changes in the UDF case and the field is expected to be exactly zero at the surface location.
@@
MeshUDF [13] NSD-UDF [36] DMUDF [41] Ours - iter. 1 Ours - iter. 2 Ours - iter. 6
@@
Most current approaches to meshing UDFs [46, 41, 14, 36] rely on Marching Cubes or Dual Contouring-like algorithms that operate on the vertices of individual voxels to recover the surface that may or may not traverse them. However neural UDFs are notoriously noisy [13, 36, 14], which makes it hard to retrieve the correct surface. Counterintuitively - and contrary to SDFs - increasing the meshing resolution to retrieve finer surface details actually worsens the problem, as shown in the results section (Fig. 5). This is because, at high resolutions, UDF values often become noisier or ambiguous within a voxel, and these local, single-pass methods lack the necessary context to infer the surface reliably, resulting in poorly recovered geometry. As shown in Fig. 1, when the neural UDF fails to reach sufficiently low values, a hole appears in the reconstructed surface. This is particularly apparent around the planes’s cockpit, as shown in the upper right corner of the figure where existing methods do not reliably detect a surface, even though there is a local minimum region. Thus, looking beyond single voxels is necessary, as attempted in [5, 13]. However, these attempts rely on basic propagation heuristics or a simple increase of the receptive fields and still fail at high resolutions, as also shown in the results section.
@@ 2 Related Work
Signed and unsigned distance fields are now in widespread use to represent 3D surfaces, which are typically parameterized by a latent vector that is decoded by a deep network [23, 27, 6]. Shapes are not tied to a specific resolution, but only to the representational power of the deep network in use. Neural signed distance fields (SDFs) are limited to watertight shapes but are relatively easy to mesh, whereas unsigned distance fields (UDFs) can represent open surfaces. They are used in volume rendering, shape reconstruction from multi-view and shape generation with a variety of techniques [20, 21, 19, 7], but they are much harder to mesh.
@@
Marching Cubes (MC) [22, 18] and Dual Contouring (DC) [15] are widely used to mesh signed fields by detecting sign transitions within voxels of a 3D grid. MC interpolates vertex positions along cell edges, while DC places them inside cells and optimizes their locations. Deep-learning adaptations, such as Neural Marching Cubes (NMC) [4] and Neural Dual Contouring (NDC) [5], improve on the classical methods by eliminating the need for manually defined rules or accurate field gradients, respectively. FlexiCubes [32] yields impressive performance on neural SDFs used for shape optimization while McGrids [29] improves the scalability of surface extraction at high resolution. In all methods increasing the resolution, or the number of points used, typically produces more detailed meshes at a higher computational cost. However, as they rely on sign changes, none of these methods is directly applicable to UDFs.
@@
NDF [6] uses a ball-pivoting algorithm to mitigate this problem. Unfortunately, its reliance on dense point clouds and expensive processing takes hours per mesh and makes it impractical. Methods such as MeshUDF [13] and CAP-UDF [44, 46] assign pseudo-signs to the field values by comparing gradients so that MC can be used, but they are subject to artifacts, especially on complex geometries. MeshUDF partially conditions its pseudo-signs on previously computed voxels by defining a specific voxel-exploration order, but it does so in a single pass and with a heuristic that targets simple shapes such as garments. DCUDF [14] extracts an $\epsilon$-level set using MC, refines it by minimizing a loss function, and cuts it to a single thin surface using a min-cut algorithm [2]. It produces smooth surfaces but requires much manual tuning, long processing times, while sometimes failing to properly cut the inflated surface. While the method refines the vertex locations with an optimization procedure, the mesh extraction is still carried out in a single pass. In NSD-UDF [36], a neural network predicts pseudo-signs starting from local UDF values and gradients, which can be meshed using SDF-like triangulation methods such as MC and DualMesh-UDF [41]. However, all these methods struggle at high resolutions on neural UDFs, where gradients and field values can become very noisy. This interferes with the proper generation of pseudo-signs, surface cuts, or dual vertices, depending on the specific method being used. This can yield inconsistencies and holes in the reconstructed surfaces, which our iterative approach is designed to prevent.
Since DC is a popular alternative to MC, it has also been adapted to work with UDFs. DualMesh-UDF [41] prunes grid cells via a UDF-based threshold and refines vertices using a DC-like quadratic equation. While preserving sharp features, it relies on hand-crafted filtering rules that can produce holes and require manual tuning. This tuning can be partially avoided by using [36] for pre-processing. However the approach still struggles at high resolutions when there is noise. Unsigned Neural Dual Contouring (UNDC) [5] uses a 3D CNN to predict dual vertex locations and connectivity. Training the network requires heavy data augmentation and ignores gradient information, limiting the method’s accuracy on neural UDFs [41, 36].
@@
Recursive refinement techniques for network predictions have been demonstrated across many domains [24, 28, 37, 33]. These approaches use the surrounding context to improve predictions [31, 9], proving especially effective for tasks such as delineation [34, 10], human pose estimation [25], semantic segmentation [42], and depth estimation [8]. These methods have inspired us to develop the first-ever iterative approach to UDF meshing.
@@
Several approaches have been proposed to learn less noisy UDFs. Zhou et al. [44, 46] enforced parallel level sets near surfaces to improve learning from sparse point clouds. Fainstein et al. [11] use hyperbolic scaling and second-order differentiation to reduce the noise in gradients and ensure the differentiability of the field. Zhao et al. [43] used anchor-based 3D position features to enhance UDF predictions, aiding single-view garment reconstruction. Wang et al. [39] predicted pseudo-signs from surface gradients for meshing, while Venkatesh et al. [38] leveraged surface orientation for better local geometry. However, precise UDF predictions remain challenging, as existing methods struggle to eliminate holes and discontinuities in meshed surfaces.
@@ 3 Method
where $\mathit{U}_{\mathcal{S}}$ is implemented by a deep network. $\mathbf{x}$ is a point in 3D space, and $d$ should be the distance from $\mathbf{x}$ to the closest points on the surface $\mathcal{S}$ corresponding to the $n$-dimensional latent vector $\mathbf{z}$ that controls the shape of $\mathcal{S}$. The difficulty with this formulation is that if $\mathit{U}_{\mathcal{S}}$ is not accurate, $\mathit{U}_{\mathcal{S}}(\mathbf{x},\mathbf{z})$ might not reach zero at places where it should, resulting in unwarranted holes in the surface reconstruction. Since unsigned fields are not differentiable everywhere, they pose a parametrization challenge, resulting in noisy representations [13, 14, 36]. Similarly, the gradients of $\mathit{U}_{\mathcal{S}}$, which are crucial for meshing [13, 44, 41, 36], can become very noisy when operating at high resolutions, also resulting in reconstruction failures.
@@ 3.1 From UDF to Triangulated Mesh
Given $\mathit{U}_{\mathcal{S}}$ of Eq. 1, we query it for UDF values and gradients on a regular grid. We then group the values in cubic cells and feed them to a neural network $f_{\theta}$, parametrized by $\theta$, whose task is to output a pseudo-sign configuration for the 8 corners of the input cell. As in [36], we use a fully connected feed forward network with $2^{7}=128$ outputs, representing a one-hot encoding of the possible pseudo-sign configurations, up to a complete sign symmetry in the cell.
@@
As in NSD-UDF [36], the final output of our pipeline is a pseudo-SDF, that is a UDF in which the grid cell corners have been assigned pseudo-signs so they can be triangulated using existing methods. We use Marching Cubes [18] for its ease of use or DualMesh-UDF [41] for its precision.
@@
Algorithms such as MeshUDF [13] and NSD-UDF [36] use manually-defined resolution-dependent thresholds that ignore cells that are too far from the surface and, thus, reduce the computational complexity of the algorithm. While this works well at lower resolutions, the high level of noise at higher resolutions makes this process error-prone, resulting in missing regions in the final mesh. Instead, we filter out cells whose vertices have a UDF value equal or above the clamping threshold of the neural UDF, set to be 0.1 in our experiments. In practice, this eliminates 85% of the cells at $256^{3}$ resolution, with no impact on the final result. For subsequent iterations, since the network outputs a probability distribution over the sign configurations, we can interpret the highest probability as the confidence of the network. We can then conservatively filter out additional cells whose confidence is extremely high (>0.999>0.999). As can be seen in Tab. 1(a), this substantially reduces the computational cost, making it comparable to that of the other baselines given in Tab. 1(b), without accuracy loss compared to meshing all cells.
Table 1: Filtering and meshing times. (a) Median L2 Chamfer Distance $\times 10^{-5}$ with 2M sample points (CD), Image Consistency (IC) and model inference times are reported at varying grid resolutions on an auto-decoder [27] trained on ShapeNet [3] cars. The times were measured on an NVIDIA A100 GPU. Notice that our filtering strategies put the proposed pipeline in the same range as existing methods, without impacting its accuracy.
@@
Time CAP [46] MeshUDF [13] DCUDF [14] NSD [36] UNDC [5] DMUDF [41] Ours
@@ 4.1 Datasets and baselines
We test our method on five different 3D categories, including garments from MGN [1]; cars, chairs, and planes from ShapeNet [3]; and natural scenes from the 3DScene dataset [47], using four different neural UDF architectures. For garments, cars, chairs, and planes, we train a traditional auto-decoder following [27, 36]. To showcase the capabilities of our method on downstream tasks and more precise single-shape neural architectures, we perform surface reconstruction from point clouds by training CAP-L [46] [^1] We refer to the learning architecture of CAP-UDF as CAP-L, to distinguish it from the meshing method introduced in the same paper, which we employ as a baseline. on 3D scenes and cars, and DiffUDF [11] on cars. In the supplementary material, we provide additional results using a softplus-based auto-decoder and meshing results obtained from ground-truth UDFs. We use 300 garments and the first 20 samples each for cars, chairs, and planes. As the MGN garments are simpler and exhibit less variety than the other shapes, we use a lower resolution for them. For 3D scenes, we use three scenes from the official CAP-UDF [46] repository [^2] The repository contains five scenes, but ”Copyroom” and ”Lounge” produced bad metrics across all methods due to inconsistent normalization, so they have been excluded..
We divide the baselines into two groups, those that rely on Marching Cubes (CAP-UDF [46], MeshUDF [13], DCUDF [14], NSD-UDF [36]) and those that use Dual Contouring (UNDC [5], DualMesh-UDF [41], NSD-UDF [36]). Notably, DCUDF [14] and DualMesh-UDF [41] have been shown to be sensitive to parameter choices. We report results obtained using both the default parameters and those that we manually tuned, denoted as “-T”. Due to implementation constraints, DualMesh-UDF and methods that depend on it use a resolution one pixel higher than the others, which we still include in the same tables. UNDC [5] failed at high resolution due to the method’s VRAM requirements. DCUDF [14] failed on certain shapes, yielding invalid metrics; these cases are excluded, making its results not directly comparable. Following official recommendations, we also evaluate DCUDF-T-nocut, a variant without the cutting step that produces double-layered meshes, on cars. All meshing methods are run without post-processing to ensure fair comparison.
@@ 4.2 Metrics
To compare extracted meshes to ground-truth surfaces, we compute three metrics: the L2 Mesh Chamfer Distance (CD) between the meshes, computed with 2M sample points as a two-way point-to-mesh distance; the Image Consistency (IC) [13], which is the product of IoU and the cosine similarity of normal maps rendered from eight viewpoints; and the F1 score [4, 5], also computed with 2M sample points, which measures the portion of the surface within an error threshold of $0.003$ from the ground-truth surface. We report the median of these metrics. It has been found to be more robust to outliers [36] than the mean, which we report in the supplementary for completeness along with the standard deviation.
@@
MGN* [1] ShapeNet cars [3] ShapeNet chairs [3] ShapeNet planes [3]
@@
128 CAP-UDF [46] 16.2 69.4 79.4 53.9 53.3 83.2 378 50.9 70.3 14.0 69.5 80.6
MeshUDF [13] 2.45 82.9 94.1 11.3 57.7 88.6 7.43 67.6 88.0 6.82 74.6 81.0
DCUDF [14] 13500 2.50 2.76 6580 7.74 14.7 17000 15.9 13.9 1880 27.6 24.5
DCUDF-T [14] 90.8 2.86 87.3 50.3 58.0 87.3 214 64.1 82.9 144 68.8 77.2
DCUDF-T-nocut [14] - - - 11.4 61.7 89.4 - - - - - -
NSD-UDF + MC [36] 1.34 83.9 94.7 6.79 59.6 88.5 6.11 67.8 88.1 3.82 79.0 84.8
@@
256 CAP-UDF [46] 1.66 86.8 91.8 34.0 61.2 87.6 114 70.5 82.0 5.50 83.7 85.4
MeshUDF [13] 0.958 89.7 95.0 13.6 62.3 88.6 27.8 72.9 87.3 3.47 85.8 85.2
DCUDF [14] 14400 4.76 3.71 346 52.6 78.2 3530 49.9 54.3 27.9 84.7 82.0
DCUDF-T [14] 4.63 86.8 95.4 347 52.6 78.2 3560 50.0 54.3 47.3 80.4 78.7
DCUDF-T-nocut [14] - - - 52.0 58.9 82.3 - - - - - -
NSD-UDF + MC [36] 0.808 90.0 95.2 10.2 62.0 87.9 10.9 72.3 86.2 2.91 87.3 86.0
@@
512 CAP-UDF [46] 0.872 90.6 94.6 31.8 61.7 87.5 63.9 71.7 82.0 5.94 87.5 86.2
MeshUDF [13] 0.798 90.6 94.8 82.7 57.0 81.7 378 61.5 65.7 12.6 88.1 84.6
DCUDF [14] 4.37 88.3 91.1 223 56.5 84.2 2950 55.0 70.1 48.7 85.5 82.2
DCUDF-T [14] 4.38 88.2 91.1 223 56.5 84.2 2000 55.0 70.1 63.0 85.4 81.3
DCUDF-T-nocut [14] - - - 43.1 60.4 85.6 - - - - - -
NSD-UDF + MC [36] 0.784 90.8 94.8 56.9 58.8 83.8 295 64.7 75.7 10.0 89.4 85.1
@@
MGN* [1] ShapeNet cars [3] ShapeNet chairs [3] ShapeNet planes [3]
@@
128 UNDC [5] 1.09 87.1 94.1 13.5 61.7 86.4 29.9 69.4 81.9 2.50 82.0 86.1
DualMesh-UDF [41] 216 68.1 68.4 952 34.4 45.5 7930 12.1 9.08 112 74.3 76.1
DualMesh-UDF-T [41] 0.806 89.9 95.4 5.56 63.5 89.5 5.34 75.1 89.1 1.96 84.3 87.5
NSD-UDF + DualMesh-UDF [36] 0.760 90.4 95.0 6.34 65.7 89.1 5.50 72.8 89.2 2.08 87.1 86.6
@@
256 UNDC [5] 0.931 89.1 91.5 82.4 52.3 71.4 293 54.2 57.8 11.6 83.6 80.7
DualMesh-UDF [41] 176 66.3 66.4 846 34.0 45.1 8280 12.4 9.26 105 77.9 76.0
DualMesh-UDF-T [41] 0.722 91.2 95.1 10.6 64.3 87.0 22.8 72.2 84.1 2.43 88.2 87.0
NSD-UDF + DualMesh-UDF [36] 0.671 91.0 94.6 10.5 64.9 87.2 14.6 70.7 84.4 2.78 88.9 85.6
@@
512 UNDC [5] 2.39 84.8 82.8 - - - - - - - - -
DualMesh-UDF [41] 167 63.7 63.9 870 32.5 43.0 8190 11.8 9.07 111 77.6 74.2
DualMesh-UDF-T [41] 0.827 90.2 93.2 37.8 58.2 79.3 72.7 63.7 72.5 4.77 89.0 84.0
NSD-UDF + DualMesh-UDF [36] 0.787 89.7 92.5 60.5 57.5 79.9 296 62.5 68.0 9.95 87.6 83.3
@@
CAP-L scenes [46] CAP-L cars [46] DiffUDF cars [11]